# Exploring the test of the no-hair theorem with LISA

### v1

@@ VI Conclusions
The advantage of the agnostic approach is that it does not require any hypothesis on the events, except for the number of QNMs (which could also be inferred by performing a Bayesian model comparison not demonstrated here). We estimate the frequency and damping time for each QNM. By comparing the mass and spin derived from the complex frequencies we identify different values for each QNM, resulting in inconsistency between QNMs in the GR framework. We also quantify these deviations as fractional deviations from GR, which entails a delicate interpretation of the results depending on the assumptions made. Indeed, the non-GR-deviation hypothesis in the dominant mode is too restrictive to correctly identify the injected deviation for each QNM, despite being consistent within $2\sigma$ with the true values. However, this kind of discrepancy can be circumvented by contrasting the results with the posterior distributions obtained from the IMR analysis, presuming that physical effects like eccentricity or others are included to avoid biased parameters. Hence this procedure requires an unbiased IMR analysis to compare results to.
The deviation approach shows better results for the fractional deviation values. However, prior assumptions are required to recover the injected values. Particularly, we assume a fixed number of observable QNMs and further constraints in the priors volume. Using the fact that no significant deviation in the first mode is observed, we do not need to rely on an IMR analysis since the intrinsic BH parameters are estimated. Including the mass and spin in the parameter estimation enables us to absorb small variations that correspond to relatively large errors in the QNMs. Thus, the hypothesis of non-deviation in the dominant mode allows us to find the injected QNM deviations confidently. If one allows for deviations in the dominant mode, extra care or further constraints in the priors are necessary due to the degeneracy between $M_{f},a_{f}$ and $\delta\omega_{220},\delta\tau_{220}$. Such an analysis is left for the future.
@@
Finally, we also evaluate the impact of redshift and total mass on the observable deviations from GR in the BH’s spectrum with the deviation template. From this analysis, we can estimate that in the best-case scenario, i.e., with “golden” sources, the strong regime of GR could be tested up to $5\times 10^{-3}\%$. However, these sources do not dominate the estimated population of black holes in the heavy or light seeds models. Nevertheless, the prospects of testing GR in the ringdown signal from sources with masses $[10^{6}-10^{7}]\,M_{\odot}$ at redshift $\leq 5$ are very promising, with a detectable fractional deviation of $\delta\omega_{330}\leq 5\times 10^{-2}\%$ in the $(3,3,0)$ mode’s frequency.
@@ Acknowledgements.
We wish to thank Sylvain Marsat for his valuable insights. This work was supported by CNES, focused on LISA Mission. We gratefully acknowledge support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing and data-processing resources needed for this work. CP acknowledges financial support from CNES Grant No. 51/20082 and CEA Grant No. 2022-039.

### v2

@@ VI Conclusions
The advantage of the agnostic approach is that it does not require any hypothesis on the events, except for the number of QNMs (which could also be inferred by performing a Bayesian model comparison not demonstrated here). We estimate the frequency and damping time for each QNM. By comparing the mass and spin derived from the complex frequencies we identify different values for each QNM, resulting in inconsistency between QNMs in the GR framework. We also quantify these deviations as fractional deviations from GR, which entails a delicate interpretation of the results depending on the assumptions made. Indeed, the non-GR-deviation hypothesis in the dominant mode is too restrictive to correctly identify the injected deviation for each QNM, despite being consistent within $2\sigma$ with the true values. However, this kind of discrepancy can be circumvented by contrasting the results with the posterior distributions obtained from the IMR analysis, presuming that physical effects like eccentricity or others are included to avoid biased parameters. Hence this procedure requires an unbiased IMR analysis to compare results.
The deviations approach shows better results for the fractional deviation values. However, prior assumptions are required to recover the injected values. Particularly, we assume a fixed number of observable QNMs and further constraints in the priors volume. Using the fact that no significant deviation in the first mode is observed, we do not need to rely on an IMR analysis since the intrinsic BH parameters are estimated. Including the mass and spin in the parameter estimation enables us to absorb small variations that correspond to relatively large errors in the QNMs. Thus, the hypothesis of non-deviation in the dominant mode allows us to find the injected QNM deviations confidently. If one allows for deviations in the dominant mode, extra care or further constraints in the priors are necessary due to the degeneracy between $M_{f},a_{f}$ and $\delta\omega_{220},\delta\tau_{220}$. Such an analysis is left for the future.
@@
Finally, we also evaluate the impact of redshift and total mass on the observable deviations from GR in the BH’s spectrum with the deviations template. From this analysis, we can estimate that in the best-case scenario, i.e., with golden sources, the strong regime of GR could be tested up to $5\times 10^{-3}\%$. This value is consistent with previous analysis [57, 98]. However, these sources do not dominate the estimated population of black holes in the heavy or light seeds models. Nevertheless, the prospects of testing GR in the ringdown signal from sources with masses $[10^{6}-10^{7}]\,M_{\odot}$ at redshift $\leq 5$ are very promising, with a detectable fractional deviation of $1\times 10^{-3}\%\leq\delta\omega_{330}\leq 5\times 10^{-2}\%$ in the $(3,3,0)$ frequency’s mode.
@@ Acknowledgements.
We wish to thank Sylvain Marsat for his valuable insights, and the reviewer for carefully read this article and suggest some adjustments. This work was supported by CNES, focused on the LISA Mission. We gratefully acknowledge support from the CNRS/IN2P3 Computing Center (Lyon - France) for providing computing and data-processing resources needed for this work. CP acknowledges financial support from CNES Grant No. 51/20082 and CEA Grant No. 2022-039.